Operation Summary for March 30, 2011 - April 4, 2011 |
The Integrated luminosity for the period from 3/28/11 to 4/4/11 was 36.95 pb-1 (inverse picobarns). NuMI was off to replace its target. |
|
Wednesday March 30 The day shift began with operators putting the Tevatron (TEV) through a check out, with Operations monitoring the Recycler stash (322.6E10) and the Antiproton Source (Pbar) stack (24.9E10), with Pbar off due to Debuncher kicker (D:IKIK) repairs, with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off due to target repairs, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, the TEV checkout continued through out the shift. Operations monitored the accelerator complex. During the evening shift, Operations put the TEV into HEP shot setup at 4:50 PM. During shot setup a TEV RF station (TRF4) tripped off and caused the TEV to quench at 6:38 PM. Operations had recovered the TEV from the quench by 10:08 PM. Operations established store 8603 with an average initial luminosity (AIL) of 259E30.Thursday March 31 The midnight shift began with Operations monitoring store 8603 (170.39E30) and the stash (111.8E10), with Pbar stacking (9.6E10), and with MiniBooNE taking beam. During the midnight shift, store 8603 quenched at 12:16 AM, due to a glitch in a vertical corrector magnet (VE47). Operations put the TEV into shot setup at 4:24 AM and established store 8606 at 5:25 AM with an AIL 214.9E30. Operations reported that the Proton Source beam quality had degraded for an hour after the Neutron Therapy Facility (NTF) had turned on, which caused beam position problems in Booster; NTF turned off before the cause of the problem could be identified. Operations began sending beam to a new Meson FTBFT-932 experiment at 5:30 AM. During the day shift, Operations monitored the accelerator complex. During the evening shift, Operations disabled beam to T-932 at 5 PM. Operations terminated store 8606 at 5:19 PM. Operations established store 8607 at 6:35 PM with an AIL of 285E30. Operations reported that Pbar and MiniBooNE experienced about 8 minutes of total downtime due to miscellaneous Booster and Linac trips and one P1 transfer line trip.Friday April 1 The midnight shift began with Operations monitoring store 8607 (126.87E30) and the stash (176.8E10), with Pbar stacking (9.5E10), and with MiniBooNE taking beam. During the midnight shift, Operations established beam to T-932 at 4:20 AM. Operations terminated store 8607. During the day shift, Operations established store 8608 at 9:12 AM with an AIL of 335.29E30. Booster personnel investigated the NTF related beam position problems throughout the day. During the evening shift, a Linac RF station (LRF1) tripped off on a bad modulator regulator at 5:51 PM and was down for a couple of hours, which halted beam to T-932 a little early.Saturday April 2 The midnight shift began with Operations monitoring the store 8608 (73.7E30) and the stash (305.7E10), with Pbar stacking (17.4E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8608 at 3:44 AM and put the TEV into shot setup, but the second transfer of antiprotons remained in MI due to a low-level RF (LLRF) cycle error at 4:23 AM. MI personnel investigated. Operations put the TEV into shot setup at 7:03 AM. During the day shift, Operations established store 8610 at 8 AM with an AIL of 346.02E30. At 11:28 AM, LRF1 tripped off due to a bad switch tube; it was down for a couple of hours. During the evening shift, Operations monitored the accelerator complex. Sunday April 3 The midnight shift began with Operations monitoring store 8610 and the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8610 at 1:52 AM, put the TEV into shot setup, and established store 8611 at 3:04 AM with an AIL of 356.49E30. During the day shift, Operations monitored the accelerator complex. During the evening shift, Operations terminated store 8611 at 6:11 PM with the intention of emptying the stash and accessing the MI tunnel to search for a LCW leak, but this didn’t happen. Operations put the TEV back into shot setup, but halted the setup when they found a quadrupole power supply (C:B0Q6) that wasn’t following its reference voltage. TEV personnel investigated and traced the problem to a loose slew rate limiter card; at 10 PM, operators reseated the card, but this resulted in a quench indication that delayed the TEV turn-on and check-out. A lightning strike at 11:52 PM, caused the fast bypass (FBS) interlock loop for the TEV power supplies to trip off. Monday April 4 The midnight shift began with the TEV in a dry squeeze, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations put the TEV into shot setup and established store 8619 at 3:08 AM with an AIL of 283.16E30. Operators prepared MI for access at 3:38 AM; operators found a LCW leak on a magnet hose at location 308. Store 8619 aborted at 6:45 AM due to a TEV RF anode power supply trip. MI personnel repaired the leaking hose and MI resumed sending beam at 7:38 AM. The Plans for Monday The plans for today are to bring the TEV back on, conduct some TEV studies, and then establish the next store. |
|
More Information Comments and Suggestions |
| last modified 04/04/2011 email Fermilab |
| Security, Privacy, Legal | |